following are the demand and supply equations for a consumer good: (20)



Qd = 300,000/P (Qd = quantity demanded)



Qs = 30P (Qs = quantity supplied)



(P = Price)




Find the market clearing price and the quantity at the equilibrium. Also sketch the demand and supply curves.

1
Expert's answer
2022-01-24T11:02:24-0500

At Equilibrium,

"Q_s=Q_d"


"\therefore 30P=\frac{300000}{P}"

"30P^2= 300000"

P= 100

Use either Equation to obtain Q

"Q_s=30P"

"Q_s=30\times 100=3000"

This information can be shown by the use of a diagram as below;